Brickhouse is expected to pay a dividend of $3.45 and $2.58 over the next two years, respectively. After that, the company is expected to increase its annual dividend at 4.1 percent. What is the stock price today if the required return is 11.9 percent?

$37.46

$30.57

$35.07

$32.63

$27.49
Answer #1

Value after year 2=(D2*Growth rate)/(Required return-Growth rate)

=(2.58*1.041)/(0.119-0.041)=$34.43307692

Current price=Future dividends*PResent value of discounting factor(11.9%,time period)

=3.45/1.119+2.58/1.119^2+34.43307692/1.119^2

=$32.63(Approx).